When Calcium carbonate is made to react with phosphoric acid, it produces calcium phosphate, carbon dioxide and water.

a.) Write a balanced equation on the above reaction.

b.) Calculate for the number of grams of calcium phosphate produced from 12.5 g of calcium carbonate and 17.3 g of phosphoric acid.

c.) How many grams of calcium phosphate produced from 3.5 moles of calcium carbonate and 7 moles of phosphoric acid?
1
Expert's answer
2021-10-15T02:54:13-0400

a) 3CaCO3 + 2H3PO4 → Ca3(PO4)2 + 3CO2 + 3H2O

b) M(CaCO3) = 100.08 g/mol

n(CaCO3) "= \frac{12.5}{100.08} = 0.125\;mol"

M(H3PO4) = 97.99 g/mol

n(H3PO4) "= \frac{17.3}{97.99} = 0.176 \;mol"

According to the reaction equation for two moles of H3PO4 we need 3 moles of CaCO3. So, CaCO3 is a limiting reactant.

n(Ca3(PO4)2) "= \frac{1}{3} n(CaCO_3) = \frac{1}{3} \times 0.125 = 0.0416 \;mol"

M(Ca3(PO4)2) = 310.18 g/mol

m(Ca3(PO4)2) "= 0.0416 \times 310.18 = 12.924 \;g"

c) According to the reaction equation for two moles of H3PO4 we need 3 moles of CaCO3. So, CaCO3 is a limiting reactant.

n(Ca3(PO4)2) "= \frac{1}{3} n(CaCO_3) = \frac{1}{3} \times 3.5 = 1.166 \;mol"

m(Ca3(PO4)2) "= 1.166 \times 310.18 = 361.876 \;g"